# Behind the Magic, MERLIM: Multi-modal Evaluation Benchmark for Large Image-Language Models

### v2

@@ Abstract
Large Vision and Language Models have enabled significant advances in fully supervised and zero-shot visual tasks. These large architectures serve as the baseline to what is currently known as Instruction Tuning Large Vision and Language models (IT-LVLMs). IT-LVLMs are general-purpose multi-modal assistants whose responses are modulated by natural language instructions and visual data. Despite this versatility, IT-LVLM effectiveness in fundamental computer vision problems remains unclear, primarily due to the absence of a standardized evaluation benchmark. This paper introduces a Multi-modal Evaluation Benchmark named MERLIM, a scalable test-bed to assess the capabilities of IT-LVLMs on fundamental computer vision tasks. MERLIM contains over 300K image-question pairs and has a strong focus on detecting cross-modal “hallucination” events in IT-LVLMs. Our results bring important insights on the performance of state-of-the-art IT-LVMLs including limitations at identifying fine-grained visual concepts, object hallucinations across tasks, and biases towards the language query. Our findings also suggest that these models have weak visual grounding, but manage to make adequate guesses from global visual patterns or language biases contained in the LLM component.
Large Visual-Language Models, Benchmark, Instruction Tuning Methods
[…]
@@ 1 Introduction
Figure 1: IT-LVLM ”hidden” Hallucinations. On the left sub-figure we show the prediction of InstructBLIP  (Liu et al., 2023), despite the small size of the tennis racquet and the relatively cluttered scene, we get an (apparently) correct response including the racquet object (we highlight in green font true positives and in red font false positives). After a subtle image edit (right sub-figure), the response remains the same although there is no visual grounding for the tennis racquet. The box around the racquet object was drawn for easier visualization, but it is not part of the original visual input.
@@
Large-scale transformer networks have significantly advanced the field of Natural Language Processing (NLP) (Brown et al., 2020; Devlin et al., 2018; Liu et al., 2019; Raffel et al., 2020). Large Language Models (LLMs) have demonstrated remarkable zero-shot performance in a wide range of NLP tasks (Lin et al., 2022; Min et al., 2023), and their success has resulted in a fundamental shift for NLP research, moving from task-specific pre-training to task-agnostic representation learning. Similarly, in the image domain, vision transformers trained on large-scale image collections have achieved state-of-the-art performance on multiple fully supervised tasks (Dosovitskiy et al., 2020; Fan et al., 2021; Liang et al., 2021). The integration of a language stream has also demonstrated empirical benefits in fundamental computer vision tasks such as recognition (Alayrac et al., 2020; Jia et al., 2021; Radford et al., 2021) and detection (Chen et al., 2021; Du et al., 2022). Ultimately, these advances in computer vision and NLP fields have converged in the development of Large Vision and Language models (LVLMs) (Alayrac et al., 2022; Kim et al., 2021; Li et al., 2022a; Nichol et al., 2021; Radford et al., 2021; Yu et al., 2022), which have achieved state-of-the-art performance on several downstream visual tasks while enabling significant improvements in the zero-shot setups.
@@
Following the success of these large-scale multi-modal transformers, a new trend has emerged which involves integrating frozen LLMs and vision transformers by means of a smaller neural network (Bai et al., 2023; Dai et al., 2023; Dong et al., 2024; Huang et al., 2023b; Li et al., 2023a; Liu et al., 2023; Peng et al., 2024; Research, 2024; Tsimpoukelli et al., 2021; Zhai et al., 2022; Zhu et al., 2023). This straightforward modification only requires the training of a fusion module to interface between the frozen modality encoders. The resulting model requires significantly less computational effort to train, and enables querying image data with natural language, thus providing a contextualized natural language response through the frozen language decoder (Tsimpoukelli et al., 2021).
@@
Such models are known as Instruction Tuning Large Vision and Language models (IT-LVLM), as their zero-shot capabilities are mediated by a natural language input that instructs the model about the expected output (Huang et al., 2023b). Although IT-LVLMs are intended to be general-purpose instructional assistants, there is still no standard test-bed to assess their zero-shot effectiveness in fundamental computer vision tasks. The presence of an LLM introduces additional challenges, among them the inherited tendency to Hallucinate (Huang et al., 2023a; Zhang et al., 2023), resulting in the emergence of language outputs that can not be grounded to the visual input.
@@
In this paper we take an in depth look at the capabilities of IT-LVLMs for fundamental computer vision tasks, while assessing the impact of hallucination events in the model’s performance. Our proposal has two key elements: First, instead of presenting the model with a fixed set of possible answers, we allow the IT-LVLM to reply in natural language and then parse the open-set IT-LVLM prediction. Second, to better analyze the hallucination issue, we propose an inpainting procedure that subtly manipulates both the image and its corresponding ground-truth. This technique enables us to identify two types of hallucinations: (1) Regular Hallucinations, which are outputs that do not correspond to the image, and (2) Hidden Hallucinations, which appear to be true positives but lack visual grounding. We illustrate these two phenomena in Figure 1, where an apparently correct prediction, such as ”Tennis Racquet,” lacks visual grounding in the edited image, thereby categorizing it as a ”hidden hallucination”.
@@
We compile our image edits, corresponding ground-truths, and parsing strategy into MERLIM (Multi-modal Evaluation benchmaRk for Large Image-language Models), a novel test-bed aimed at empirically evaluating IT-LVLMs on core computer vision tasks including object recognition, instance counting and identifying object-to-object relationships. Our benchmark serves as a straightforward and extensible evaluation strategy of core computer vision tasks on well-established image datasets like MS-COCO (Lin et al., 2014), LVIS (Gupta et al., 2019), and Visual Genome.
@@
Our paper brings the following contributions: (i) A standardized test-bed to assess the zero-shot effectiveness of IT-LVLMs on fundamental computer vision tasks, including Object Recognition, Object Counting, and Inter-object Relationship Understanding. (ii) Through extensive empirical evaluation, we identify and quantify ”hidden” hallucination errors which are out of the scope of other benchmarks, we observe this hallucination unfairly increase the accuracy of IT-LVMLMs. (iii) We conducted a thorough analysis of language biases through extensive experiments using a diverse set of carefully selected question prompts. Our observations reveal that IT-LVLMs exhibit a strong bias toward the language tokens of both the question and answer.
@@ 2 Related Work
The widespread adoption of LLMs and their zero-shot capabilities (Brown et al., 2020; Chiang et al., 2023; Devlin et al., 2018; Liu et al., 2019; Raffel et al., 2020) has led to an increased interest in the development of LVLMs (Jia et al., 2021; Pham et al., 2023; Radford et al., 2021; Yu et al., 2022; Zhai et al., 2022). LVLMs typically build upon metric learning, estimating a similarity measure across modalities, making these architectures effective in inherently multi-modal tasks such as VQA and text-image retrieval.
@@
The effectiveness of LVLMs, coupled with the unprecedented availability of large-scale data in the visual domain, has led to the development of IT-LVLMs. Unlike LVLMs, IT-LVLMs allow to prompt a joint language and visual representation and then generate results with a frozen language decoder. This procedure preserves the multi-modal capabilities of the LVLM, and enables the use of natural language queries as an instruction set that modulates the expected output(Alayrac et al., 2022; Tsimpoukelli et al., 2021).
A pioneer work exploring the few-shot learning abilities of frozen language models and vision was presented by Tsimpoukelli et al. (Tsimpoukelli et al., 2021) where the image input is encoded into a feature set and inserted as a prefix in the text stream, thereby conditioning the responses of the language network. Flamingo (Alayrac et al., 2022) introduced an auto-regressive text model conditioned on visual input, by using gated cross-attention. The tokens from the perceiver architecture (Jaegle et al., 2021) are fed into a frozen language model. This Gated cross-attention was extended by BLIP-2 (Li et al., 2023a), linking the T5 language model and its visual stream (Raffel et al., 2020) with the Q-Former module, thus improving the modality alignment at training time. Recent work has continued to explore the visual-to-language alignments on IT-LVLMS, either by improving the training and annotation data (Liu et al., 2023; Zhu et al., 2023), or by introducing large-scale data on the training of the IT-LVLM (Bai et al., 2023; Dai et al., 2023).
@@
Many existing benchmarks and datasets for IT-LVLMs are designed for the standard VQA tasks, including a pre-defined set of questions that can be answered by looking at image or video data (Antol et al., 2015; Chen et al., 2015; Fu et al., 2023; Goyal et al., 2017; Gurari et al., 2018; Hudson & Manning, 2019; Zhao et al., 2022). Although these benchmarks provide a direct evaluation of the multi-modal capabilities of LVLMs, they do not assess if the reply originated from the global visual context of the image or from the local visual patterns that effectively ground the prompted question. In contrast, MERLIM focuses on querying and attributing the response of the IT-LVLM to a specific set of object instances in the image by means of small image alterations and direct verification of the same query on the modified visual data. Moreover, we explore the connection between these replies (original image and edited image) and the potential hallucinations from the LLMs component, as some responses can only be attributed to the LLM hallucinating objects.
Perhaps the most similar benchmarks to ours are POPE (Li et al., 2023b) and NExT-GQA (Xiao et al., 2023), as both these benchmarks aim at finding direct visual groundings related to the language query. While POPE proposes to prompt for the presence/absence of a single instance to evaluate hallucinations (in the form of a false positive), we directly modify the visual ground-truth and can verify if the predictions have an effective visual grounding. NExT-GQA approximates the grounding of a prediction by the accurate temporal localization of the event. Despite the straightforward proposal, contemporary research (Otani et al., 2020) has shown that temporal action localization commonly suffers from biased predictions, in particular, this bias emerges directly from the language stream.
@@ 3 MERLIM
MERLIM’s main goal is to assess the effectiveness of Instruction-Tuning Large Visual Language Models (IT-LVLM) in fundamental computer vision tasks, while simultaneously evaluating if the responses have an effective visual grounding. In practice, we design an evaluation scheme that queries the IT-LVLM with a set of predefined language questions. Then, we parse the text responses, transforming the natural language output into a suitable prediction format for each task. This scheme enables us to measure the performance of the IT-LVLM using standard image evaluation metrics, such as accuracy, precision, recall, and F1 metrics. Across all tasks, MERLIM contains a total of 300.664 unique image-question pairs. Table 1 outlines the sizes of the query set and the image data for all the selected tasks.
@@
A pre-trained IT-LVLM ($h$) takes as input an image ($i$) and a language query ($l$), producing a language response ($r$). Formally: $r=h(f(i,\theta_{f}),g(l,\theta_{g}),\theta_{h})$, where $f$ is an image encoder with parameter set $\theta_{f}$, $g$ is a language encoder with parameter set $\theta_{g}$, and $\theta_{h}$ are IT-LVLM specific parameters that integrate the two modalities. In MERLIM, we define a fixed language query set ($L$), where $l\in L$, and transform the open set output $r$ into a set $r^{\prime}\in K$ where $K$ can be directly matched to the closed-set ground truth of fundamental computer vision tasks.
@@
Object Recognition Set/Enumeration 5 4.183 31.373
Object Counting Numeric 2 4.183 31.373
Object Relationship Underst. True / False 4 1.240 5.630
Table 1: MERLIM Benchmark. We incorporate three core visual tasks to benchmark IT-LVLMs: Object Recognition, Object Counting, and Relationship Understanding. We leverage images from the validation set of MS-COCO (Lin et al., 2014) and create edited versions of these images (”Edited” column). Moreover, we extend the MS-COCO ground-truth classes by incorporating the class labels present in LVIS (Gupta et al., 2019) and include the relationships proposed in the overlapping set of Visual Genome (Krishna et al., 2017).
@@
We focus our study on three core computer vision tasks: (i) Object Recognition, (ii) Object Relationship Understanding, and (iii) Object Counting. The primary source of data for our Benchmark is the validation set of MS-COCO, where we can directly evaluate the Object recognition and instance counting. For the Reasoning Task, we resort to the labels provided on the validation set of Visual Genome (Krishna et al., 2017).
Considering the open nature of the vocabulary on IT-LVLMs responses, some correctly grounded visual objects could be labeled as false predictions if their class name is not included in the $80$ Categories of MS-COCO ground truth labels. To alleviate this undesired penalty, we complement the MS-COCO ground-truths with the labels provided by the LVIS dataset (Gupta et al., 2019). This extended ground-truth allows us to compare against a richer vocabulary.
@@
In MERLIM, we assess the effective visual grounding of the IT-LVLM responses through two separate tests. First, we directly compare the responses against the dataset ground-truth to identify Regular Hallucinations, which occur when the predicted noun is not present in the image ground truth. Second, we edit images such that one of the object instances disappears and then compare the IT-LVLM responses when presented with the original ($i$) and edited ($\hat{i}$) versions. This allows us to identify Hidden Hallucinations, which correspond to correct predictions that lack any visual grounding. These errors are beyond the scope of any other multi-modal hallucination benchmarks like (Li et al., 2023b; Xiao et al., 2023). Figure 1 illustrates this process: the original image (left) includes a tennis racquet (near the leg of the child located on the far left), while in the edited version (right), the tennis racket was inpainted. Despite the absence of the racquet, InstructBLIP with Vicuna7B continues to infer its presence. To minimize the alterations to the original image, we remove only one object instance at a time. We use the inpainting method proposed by Li et al. (Li et al., 2022b) to fill in the ground-truth segmentation mask according to the image context. The edited validation set expands to a total of $31373$ inpainted images obtained from the $4183$ original images in MS-COCO val. We empirically verify that our edits do not significantly alter the global or local appearance of the image ($\hat{i}$), we include a detailed analysis of this evaluation in the supplementary material.

### v3

@@ Abstract
Large Vision and Language Models have enabled significant advances in fully supervised and zero-shot visual tasks. These large architectures serve as the baseline to what is currently known as Instruction Tuning Large Vision and Language models (IT-LVLMs). IT-LVLMs are general-purpose multi-modal assistants whose responses are modulated by natural language instructions and visual data. Despite this versatility, IT-LVLM effectiveness in fundamental computer vision problems remains unclear, primarily due to the absence of a standardized evaluation benchmark. This paper introduces a Multi-modal Evaluation Benchmark named MERLIM, a scalable test-bed to assess the capabilities of IT-LVLMs on fundamental computer vision tasks. MERLIM contains over 300K image-question pairs and has a strong focus on detecting cross-modal “hallucination” events in IT-LVLMs. Our results bring important insights on the performance of state-of-the-art IT-LVLMs including limitations at identifying fine-grained visual concepts, object hallucinations across tasks, and biases towards the language query. Our findings also suggest that these models have weak visual grounding, but manage to make adequate guesses from global visual patterns or language biases contained in the LLM component. We name this phenomenon of correct answers with no visual grounding as hidden hallucinations.
@@ 1 Introduction
Large-scale transformer networks have significantly advanced the field of Natural Language Processing (NLP) [5, 10, 37, 46]. Large Language Models (LLMs) have demonstrated remarkable zero-shot performance in a wide range of NLP tasks [34, 39], and their success has resulted in a fundamental shift for NLP research, moving from task-specific pre-training to task-agnostic representation learning. Similarly, in the image domain, vision transformers trained on large-scale image collections have achieved state-of-the-art performance on multiple fully supervised tasks [12, 14, 33]. The integration of a language stream has also demonstrated empirical benefits in fundamental computer vision tasks such as recognition [1, 26, 45] and detection [6, 13]. Ultimately, these advances in computer vision and NLP fields have converged in the development of Large Vision and Language models (LVLMs) [2, 27, 29, 40, 45, 53], which have achieved state-of-the-art performance on several downstream visual tasks while enabling significant improvements in the zero-shot setups.
@@
Following the success of these large-scale multi-modal transformers, a new trend has emerged which involves integrating frozen LLMs and vision transformers by means of a smaller neural network [4, 9, 11, 23, 30, 36, 43, 52, 49, 54, 58]. This straightforward modification only requires the training of a fusion module to interface between the frozen modality encoders. The resulting model requires significantly less computational effort to train, and enables querying image data with natural language, thus providing a contextualized natural language response through the frozen language decoder [49].
@@
Such models are known as Instruction Tuning Large Vision and Language models (IT-LVLM), as their zero-shot capabilities are mediated by a natural language input that instructs the model about the expected output [23]. Although IT-LVLMs are intended to be general-purpose instructional assistants, there is still no standard test-bed to assess their zero-shot effectiveness in fundamental computer vision tasks. The presence of an LLM introduces additional challenges, among them the inherited tendency to Hallucinate [22, 56], resulting in the emergence of language outputs that can not be grounded to the visual input.
@@
In this paper we take an in depth look at the capabilities of IT-LVLMs for fundamental computer vision tasks, while assessing the impact of hallucination events in the model’s performance. Our proposal has two key elements: First, instead of presenting the model with a fixed set of possible answers, we allow the IT-LVLM to reply in natural language and then parse the open-set IT-LVLM prediction. Second, to better analyze the hallucination issue, we propose an inpainting procedure that subtly manipulates both the image and its corresponding ground-truth. This technique enables us to identify two types of hallucinations: (1) Regular Hallucinations, which are outputs that do not correspond to the image, and (2) Hidden Hallucinations, which appear to be true positives but lack visual grounding. We illustrate these two phenomena in Figure 1, where an ostensibly correct prediction, such as "Tennis Racquet," lacks visual grounding in the edited image, thereby categorizing it as a "hidden hallucination".
@@
We compile our image edits, corresponding ground-truths, and parsing strategy into MERLIM (Multi-modal Evaluation benchmaRk for Large Image-language Models), a novel test-bed aimed at empirically evaluating IT-LVLMs on core computer vision tasks including object recognition, instance counting and identifying object-to-object relationships. Our benchmark serves as a straightforward and extensible evaluation strategy of core computer vision tasks on well-established image datasets like MS-COCO [35], LVIS [17], and Visual Genome.
@@
Our paper brings the following contributions: (i) A standardized test-bed to assess the zero-shot effectiveness of IT-LVLMs on fundamental computer vision tasks, including Object Recognition, Object Counting, and Inter-object Relationship Understanding. (ii) Through extensive empirical evaluation, we identify and quantify "hidden" hallucination errors which are out of the scope of other benchmarks, we observe this hallucination unfairly increase the accuracy of IT-LVMLMs. (iii) We conducted a thorough analysis of language biases through extensive experiments using a diverse set of carefully selected question prompts. Our observations reveal that IT-LVLMs exhibit a strong bias toward the language tokens of both the question and answer. MERLIM’s code is available here
Figure 1: IT-LVLM "hidden" Hallucinations. On the left sub-figure we show the prediction of InstructBLIP  [36], despite the small size of the tennis racquet and the relatively cluttered scene, we get an (apparently) correct response including the racquet object (we highlight in green font true positives and in red font false positives). After a subtle image edit (right sub-figure), the response remains the same although there is no visual grounding for the tennis racquet. The box around the racquet was drawn for easier visualization, but it is not part of the original visual input.
@@ 2 Related Work
The widespread adoption of LLMs and their zero-shot capabilities [5, 8, 10, 37, 46] has led to an increased interest in the development of LVLMs [26, 44, 45, 53, 54]. LVLMs typically build upon metric learning, estimating a similarity measure across modalities, making these architectures effective in inherently multi-modal tasks such as VQA and text-image retrieval.
@@
The effectiveness of LVLMs, coupled with the unprecedented availability of large-scale data in the visual domain, has led to the development of IT-LVLMs. Unlike LVLMs, IT-LVLMs allow to prompt a joint language and visual representation and then generate results with a frozen language decoder. This procedure preserves the multi-modal capabilities of the LVLM, and enables the use of natural language queries as an instruction set that modulates the expected output [2, 49].
A pioneer work exploring the few-shot learning abilities of frozen language models and vision was presented by [49] where the image input is encoded into a feature set and inserted as a prefix in the text stream, thereby conditioning the responses of the language network. Flamingo [2] introduced an auto-regressive text model conditioned on visual input, by using gated cross-attention. The tokens from the perceiver architecture [25] are fed into a frozen language model. This Gated cross-attention was extended by BLIP-2 [30], linking the T5 language model and its visual stream [46] with the Q-Former module, thus improving the modality alignment at training time. Recent work has continued to explore the visual-to-language alignments on IT-LVLMS, either by improving the training and annotation data [36, 58], or by introducing large-scale data on the training of the IT-LVLM [4, 9].
@@
Many existing benchmarks and datasets for IT-LVLMs are designed for the standard VQA tasks, including a pre-defined set of questions that can be answered by looking at image or video data [3, 7, 15, 16, 18, 24, 57]. Although these benchmarks provide a direct evaluation of the multi-modal capabilities of LVLMs, they do not assess if the reply originated from the global visual context of the image or from the local visual patterns that effectively ground the prompted question. In contrast, MERLIM focuses on querying and attributing the response of the IT-LVLM to a specific set of object instances in the image by means of small image alterations and direct verification of the same query on the modified visual data. Moreover, we explore the connection between these replies (original image and edited image) and the potential hallucinations from the LLMs component, as some responses can only be attributed to the LLM hallucinating objects.
Perhaps the most similar benchmarks to ours are POPE [32] and NExT-GQA [51], as both these benchmarks aim at finding direct visual groundings related to the language query. While POPE proposes to prompt for the presence/absence of a single instance to evaluate hallucinations (in the form of a false positive), we directly modify the visual ground-truth and can verify if the predictions have an effective visual grounding. NExT-GQA approximates the grounding of a prediction by the accurate temporal localization of the event. Despite the straightforward proposal, contemporary research [42] has shown that temporal action localization commonly suffers from biased predictions, in particular, this bias emerges directly from the language stream.
@@ 3 MERLIM
MERLIM’s main goal is to assess the effectiveness of Instruction-Tuning Large Visual Language Models (IT-LVLM) in fundamental computer vision tasks, while simultaneously evaluating if the responses have an effective visual grounding. In practice, we design an evaluation scheme that queries the IT-LVLM with a set of predefined language questions. Then, we parse the text responses, transforming the natural language output into a suitable prediction format for each task. This scheme enables us to measure the performance of the IT-LVLM using standard image evaluation metrics, such as accuracy, precision, recall, and F1 metrics. Across all tasks, MERLIM contains a total of 300,664 unique image-question pairs. Table 1 outlines the sizes of the query set and the image data for all the selected tasks.
@@
A pre-trained IT-LVLM ($h$) takes as input an image ($i$) and a language query ($l$), producing a language response ($r$). Formally: $r=h(f(i,\theta_{f}),g(l,\theta_{g}),\theta_{h})$, where $f$ is an image encoder with parameters $\theta_{f}$, $g$ is a language encoder with parameters $\theta_{g}$, and $\theta_{h}$ are IT-LVLM parameters that connect the two modalities. In MERLIM, we define a fixed language query set ($L$), where $l\in L$, and transform the open set output $r$ into a set $r^{\prime}\in K$ where $K$ can be directly matched to the closed-set ground truth of fundamental computer vision tasks.
Table 1: MERLIM Benchmark. We incorporate three core visual tasks to benchmark IT-LVLMs: Object Recognition, Object Counting, and Relationship Understanding. We leverage images from the validation set of MS-COCO [35] and create edited versions of these images ("Edited" column). Moreover, we extend the MS-COCO ground-truth classes by incorporating the class labels present in LVIS [17] and include the relationships proposed in the overlapping set of Visual Genome [28].
@@
Object Recognition Set/Enumeration 5 4183 31373
Object Counting Numeric 2 4183 31373
Relationship Underst. True / False 4 1240 5630
@@
We focus our study on three core computer vision tasks: (i) Object Recognition, (ii) Object Relationship Understanding, and (iii) Object Counting. The primary source of data for our Benchmark is the validation set of MS-COCO, where we can directly evaluate the Object recognition and instance counting. For the Reasoning Task, we resort to the labels provided on the validation set of Visual Genome [28].
Considering the open nature of the vocabulary on IT-LVLMs responses, some correctly grounded visual objects could be labeled as false predictions if their class name is not included in the $80$ Categories of MS-COCO ground truth labels. To alleviate this undesired penalty, we complement the MS-COCO ground truths with the labels provided by the LVIS dataset [17]. This extended ground truth allows us to compare against a richer vocabulary.
@@
In MERLIM, we assess the effective visual grounding of the IT-LVLM responses through two separate tests. First, we directly compare the responses against the dataset ground truth to identify Regular Hallucinations, which occur when the predicted noun is not present in the image ground truth. Second, we edit images such that one of the object instances disappears and then compare the IT-LVLM responses when presented with the original ($i$) and edited ($\hat{i}$) versions. This allows us to identify Hidden Hallucinations, which correspond to correct predictions that lack any visual grounding. These errors are beyond the scope of any other multi-modal hallucination benchmarks like [32, 51]. Figure 1 illustrates this process: the original image (left) includes a tennis racquet (near the leg of the child located on the far left), while in the edited version (right), the tennis racket was inpainted. Despite the absence of the racquet, InstructBLIP with Vicuna7B continues to infer its presence. To minimize the alterations to the original image, we remove only one object instance at a time. We use the inpainting method proposed by [31] to fill in the ground-truth segmentation mask according to the image context. The edited validation set expands to a total of $31373$ inpainted images obtained from the $4183$ original images in MS-COCO val. We empirically verify that our edits do not significantly alter the global or local appearance of the image ($\hat{i}$), we include a detailed analysis of this evaluation in the Supplementary Material.